OpenStudy (anonymous):

Evaluate m = y2-y1/x2-x1 for for m = 1 and (x1, y1) = (0, 2) and (x2, y2) = (x, 0).

OpenStudy (anonymous):

if m = (y2 - y1)/(x2 - x1) =(0 - 2)/(x-0) => m = -(2/x) => 1 = -(2/x) {as m = 1} => x = -2